A truck starts out at position x = 3.1 mi and y = 6.5 mi and winds up at position x = 4.6 mi, y = 2.1 mi. Find the magnitude and direction of the truck's displacement.

magnitude =  mi
direction =  
Answer #1

displacement = 4.6i + 2.1j - (3.1i + 6.5j) = 1.5i-4.4j

magnitude = [1.5^2 + 4.4^2]^0.5 = 4.648 miles

direction = tan^-1 (-4.4/1.5) = -71.175 degrees or 288.825 degrees counterclockwise